Benoit, Inc., a manufacturer of plastic products, reports the following manufacturing costs and account analysis classificatiAccount Direct materials Direct manufacturing labour Power Supervision labour Materials-handling labour Maintenance labour DeBenoit, Inc., produced 80,000 units of product in 2017. Benoits management is estimating costs for 2018 on the basis of 2017a. Direct materials prices in 2018 are expected to increase by 8% compared with 2017 b. Under the terms of the labour contrac1. Prepare a schedule of variable, fixed, and total manufacturing costs for each account category in 2018. Estimate total manRequirement 1. Prepare a schedule of variable, fixed, and total manufacturing costs for each account category in 2018. Estimain 2018 Account Direct materials Direct manufacturing labour Power Supervision labour Materials-handling labour Maintenance l
Answer #1

Table is showing per unit cost for 2017 with 80000 units and 2018 with 85000 units -  

Particulars Per unit cost

2017

(80000 units)

Per unit cost

2018

(85000 units)

Direct Materials 4 320000 4.32 367200
Direct Manufacturing Cost 3 240000 3,36 285600
Power 0.50 40000 0.47 40000

Supervision labour - ( 68000/ 80000= 0.85)

Variable( 20%) ( 0.85 X 20%)

Fixed ( 80%)

0.17

0.68

13600

54400

0.17

0.64

14450

54400

Materials Handling labour - (60000/80000=0.75)

Variable (40%)

Fixed (60%)

0.30

0.45

24000

36000

0.30

0.42

25500

36000

Maintenance Labour - ( 80000/80000=1)

Variable (45%)

Fixed (55%)

0.45

0.55

36000

44000

0.45

0.52

38250

44000

Depreciation (Fixed) 1.125 90000 1.11 94500
Rent, Property taxes and administration (Fixed) 1.56 125000 1,62 137500
Total 12.79 1023000 13.38 1137400

Total manufacturing cost per unit in 2017 is 12.79 for 80000 units.

   Total manufacturing cost per unit in 2018 is 13.38 for 85000 units.

This is very better estimates to fixed and variable cost. And this estimated is helpful to Benoit plastic manufacturing. and its showing that if prices are increasing then increased in production is showing increasing per unit cost.so this is helpful estimate.